### v1

# Combinatorial Auctions without a Numeraire: The Case of Blockchain Trade-Intent Auctions

Abstract
Blockchain trade intent auctions currently intermediate approximately USD 5 billion monthly. Due to production complementarities, the auction is combinatorial: when multiple trade intents from different traders are auctioned off simultaneously, a bidder (here called solver) can generate additional efficiencies by winning a batch of multiple trade intents. However, unlike other combinatorial auctions studied in the literature, the auction has no numeraire. Fairness is a concern as the efficiencies from batching cannot be easily shared between traders. We formalize this problem and study the most commonly used auction formats: batch auctions and multiple simultaneous auctions. We also propose a novel fair combinatorial auction that combines batch auction and multiple simultaneous auctions: solvers submit individual-trade bids and batched bids, but batched bids are considered only if they are better for all traders relative to the outcome of multiple simultaneous auctions (constructed using the individual-trade bids). We find a trade-off between the fairness guarantees provided by the auction (i.e., the minimum each trader can expect to receive) and the expected value of the assets returned to the traders. Also, the amount that each trader receives in the equilibrium of the fair combinatorial auction may be higher or lower than what they receive in the equilibrium of the simultaneous auctions used as a benchmark for fairness.
August 22, 2024  (first draft, all comments are welcomed)
1 Introduction
Permissionless blockchains like Ethereum allow anyone to create blockchain-based applications at minimal or no cost, including blockchain-based financial markets. As a result, there are now dozens of such markets, each with a different mechanism, fee structure, and liquidity. The resulting market fragmentation generates complexity for users because executing a trade most efficiently may require accessing multiple marketplaces, often via intermediate assets. Also, these complex executions expose users to “sandwich attack”, in which an attacker front-runs a victim’s trade with a trade in the same direction and then back-runs it with a trade in the opposite direction. Doing so allows the attacker to “buy cheap” and “sell expensive” while forcing the victim to trade at less favorable terms.
Trade-intent auctions emerged as a response to this complexity and these risks. A trade intent is a “partial blockchain transaction” that specifies an amount of a sell asset and a buy asset (or an amount of a buy asset and a sell asset) without specifying how the transaction should be executed. The right to complete this partial transaction is auctioned off to specialized entities called solvers, who compete to provide the best possible execution to each trade intent. The first notable trade-intent auction was CoW Swap’s batch auction, in which a group of trade intents are actioned off simultaneously, solvers submit a single bid that may contain multiple trade intents, and the auctioneer selects a single winner. Later entrants, such as 1inch (1inch fusion) and Uniswap (UniswapX), instead hold Dutch or descending price auctions, one per each trade intent. Overall, trade-intent auctions currently intermediate approximately USD 5 billion monthly. [^1] More precisely, the total volume of trade-intent auctions on Ethereum was USD 5.9 billion in April 2024, USD 7.7 billion in March 2024, USD 4.9 billion in May 2024, USD 4.2 billion in June 2024, USD 5.2 billion in July 2024 (source: https://dune.com/queries/3058985/5091785). For comparison, the widely studied UK 3G spectrum auction held in 2001 generated USD 34 billion (or approx USD 60 billion at current prices).
From the solvers’ perspective, competing in the auction requires a balancing act between specialization in certain trade intents and the ability to exploit complementarities between trade intents. To specialize, some solvers may develop sophisticated algorithms to access public marketplaces, while others may work closely with private market makers. To exploit complementarities between trade intents, solvers may find coincidences of wants allowing traders to trade with each other, sometimes in complex ring trades or intermediate steps of multi-step trades. Additionally, because blockchain fees (i.e., gas costs) are largely independent of trade volume, two orders may save on fees if executed together. Solvers also face the risk of sandwich attacks and employ complex strategies to mitigate them.
From the theoretical viewpoint, trade-intent auctions are combinatorial because how a solver can best execute a trade depends on the other trades the solver also executes. Similarly to most combinatorial auctions, the goal of the auction designer is to balance the benefit of specialization (i.e., allocating each trade intent to the solver who is best at that trade intent) with the additional efficiencies from batching (i.e., allocating multiple orders to the same solver). But unlike most combinatorial auctions, the assets exchanged in the auctions are illiquid: the premise of the auction is that even if each asset has a notional market value, exchanging them is subject to frictions and fees. Hence, combinatorial auctions lack a common numeraire to reallocate the efficiencies from batching between traders.
As an illustrative example, consider two traders (who may live anywhere in the world), one demanding DOGE (a cryptocurrency with a total market capitalization of USD 10B) and the other demanding ETH (the native currency of the Ethereum blockchain, with a current market capitalization of USD 250B). Two solvers submit a bid each for the entire batch (i.e., for both trades). Suppose that the bid from the first solver generates more DOGE, fewer ETH, and has a higher overall market value than the second solver’s bid. Although each trader probably values receiving DOGE, it is unlikely that they value it in the same way — that is, it is unlikely that the auctioneer can transfer utility between the traders seamlessly by allocating the extra DOGE between them. Alternatively, the auctioneer or the solvers could trade DOGE for ETH. However, due to market frictions and fees, the trader demanding ETH may prefer the bid by the second solver even after such an exchange. As a consequence, the two traders may have opposite preferences over how to allocate trades to solvers.
This paper proposes a theoretical model of trade intent auctions. In the model, two traders demand two different assets, and two solvers can produce these assets. There are benefits from specialization because each solver has an absolute advantage in a given trade. There are also complementarities because each solver can generate more of the asset demanded by a trader when this solver also has access to the other trade. The main assumption is that there are notional market prices for the assets demanded, and it is always possible to evaluate the solvers’ bids (whether on individual trades or both trades) by their dollar value. However, the assets demanded are illiquid within the auction; that is, exchanging them for each other or dollars before the end of the auction is subject to frictions and fees. Formally, we introduce a novel feasibility constraint imposing an upper bound to the amount of each asset that solvers can return, where this upper bound is a function of the allocation of trades to solvers.
We then use the model to compare the most commonly used auction mechanisms: batch action and simultaneous standard auctions (as in Gentry et al., 2019). At a superficial level, one could think that simultaneous standard auctions are better at exploiting the benefit of specialization because different solvers may win different orders. Batch auctions, instead, are better at exploiting the complementarities between various orders. Hence, simultaneous standard auctions should be preferred when the benefit of specialization is large, while batch auctions should be preferred when the complementarities between different orders are strong. This reasoning, however, misses an important element: when the benefit of specialization is strong, competition is low in the simultaneous standard auctions. In other words, specialization implies that, in each standard auction, there is a “leader” who can win the auction easily — that is, can win while returning little to the traders. On the other hand, batching forces all solvers to compete with each other: each “leader” in a given type of trade needs to outbid the other leaders to win any trade. Hence, solvers have a higher incentive to bid (and return a higher fraction of the assets produced) in the batch auction relative to the simultaneous standard auctions. At the same time, the batch auction has multiple equilibria because the solvers’ payoff only depends on the value of the assets returned to the traders, not how this value is shared. Therefore, the benefit of the batch auction may accrue disproportionally to one trader, leaving the other worse off relative to the simultaneous standard auctions. At an intuitive level, such an outcome seems unfair. [^2] CoW Swap implementation of the batch auction tackles this problem by adding a period to the game: after the assets are returned to the users, CoW Swap checks the solver’s execution against the prices on the most popular blockchain-based financial marketplaces. If the solver’s execution of an order falls below such reference, CoW Swap asks the solver to reimburse the user. The main drawback is that the benchmark to define whether the execution of an order is fair is determined by CoW Swap and does not incorporate solvers’ knowledge.
Motivated by this observation, we then introduce a definition of fairness inspired by the literature on cooperative game theory. A weak notion of fairness is that traders are batched together if and only if they all benefit relative to a reference outcome. In cooperative game theory, this reference outcome is the players’ “outside options”: the highest possible payoff each player can achieve without the collaboration of the other players. In the case of the combinatorial auction, instead, we assume that the reference outcome is the auction’s outcome when the batched bids are removed. [^3] If multiple batched bids are better than the reference outcome for both players, then we can strengthen the notion of fairness, for example, by ranking different allocations by the product of the extra utility relative to the benchmark. As it will become clear later, our main results continue to hold for any strengthening of the notion of fairness. The reference outcome is, therefore, constructed as part of the mechanism and depends on the solvers’ bids. For ease of comparison with the previously discussed mechanisms, when solving for the equilibrium of the fair combinatorial auction, we assume that the auction can be decomposed into two individual trade auctions, which can be in first price or second price, and a batched auction in which the winner “pays its bid” (i.e., it returns to the traders the amounts specified in its batched bid). Fairness implies that the auctioneer ignores a batched bid whenever it delivers less than the outcome of the two individual trade auctions, and it always chooses a batched bid whenever there is one that is fair.
The fairness assumption introduces two strategic considerations. First, a solver’s bids in the individual trade auctions affect what batched bids are considered fair. Hence, relative to two simultaneous standard auctions, solvers may bid higher in the individual trade auctions of the fair combinatorial auction to exclude the opponent’s batched bid as unfair. The second effect is a consequence of the first one: a solver who anticipates that the opponent’s batched bid will be excluded because unfair also anticipates lower competition for the entire batch. Hence, a solver may decrease the batched bid relative to the batched auction. The strength of these effects depends on the specific auction format, that is, whether the individual-trade auctions are in first or second price.
If the individual-trade auctions are in second price, then each solver has limited ability to influence the fairness benchmark: a solver who bids higher than the opponent in the individual orders can never increase the fairness benchmark, and a solver who bids lower than the opponent can increase the fairness benchmark just up to the opponent’s bids. As a consequence, in equilibrium, the auction may not provide any fairness guarantee. For example, if both solvers set their individual-trade bids to zero (which is equivalent to not bidding on individual trade), then the equilibrium is always batching, and the only additional “fairness” guarantee provided by auction is that each trader should receive a strictly positive amount of the asset demanded. In practice, the two effects described earlier are mute, and the outcome of this fair combinatorial auction is identical to that of the batched auction.
If, instead, the individual-trade auctions are in first price, then each solver can manipulate the fairness benchmark with his individual-trade bids. The two effects described earlier are present, and solvers will bid higher in the individual-trade auction but lower in the batched bid. In equilibrium, we find three possible regimes:
  • Specialization: if, on each trade, the difference in solvers’ productivity is very large, then each solver will disqualify the opponent’s batched bid as unfair. The outcome is specialization, and each trader receives the assets that the worse solver at his order would provide when matched with both orders. This last part is important because it implies that traders receive more than what they would receive in a simultaneous standard auction (as those held by 1 Inch and Uniswap).
  • Competitive Batching: when the benefit of batching is large relative to the traders’ productivity, then the equilibrium is similar to that of the batch auction, with one important difference: the equilibrium is unique, and each trader receives at least the assets that the worst solver at his order would provide when matched with both orders.
  • Uncompetitive Batching: if the benefit of batching is in an intermediate range, and one of the two solvers is particularly productive relative to the other, then this solver can easily disqualify the opponent’s bid as unfair and hence win both traders with a relatively low bid. Like under specialization, in equilibrium, each trader receives the assets that the worst solver at his order would provide when matched with both orders.
Hence, relative to the fair combinatorial auction in which the individual-trade auctions are in the second price, when the individual-trade auctions are in first price two things happen. First, the auction provides stronger fairness guarantees. At the same time, we show that the market value of the assets delivered to the users is lower, strictly so if the outcome is specialization or uncompetitive batching.
The comparison between the two mechanisms, therefore, illustrates two important points. First, in equilibrium, higher fairness guarantees come at the expense of the market value of the assets delivered. Second, the minimum that each trader receives in equilibrium may be very different from the equilibrium of the mechanism used as a benchmark for fairness. For example, when the mechanism used as a reference for fairness is two second-price auctions, one of the traders may earn less in the equilibrium of the fair combinatorial auction than in the equilibrium of the second-price auctions. If the reference for fairness is two first-price auctions, then both traders always earn more in the equilibrium of the fair combinatorial auction than in the equilibrium of the first-price auctions.
The remainder of the paper proceeds as follows. Next, we discuss the relevant literature. Then, we introduce the model. In the following section, we use the model to study the two commonly used auction formats for trade intents. In the following section, we introduce our notion of fairness. Then, we derive the equilibrium of the fair combinatorial auction. Section 6 discusses some extensions to the model. The last section concludes. Lengthy mathematical derivations missing from the text are relegated to the appendix.
1.1 Relevant literature
This paper is related to several strands of literature. Che and Gale (1998) study a single auction with financially constrained bidders and show that a first-price auction generates higher revenues than a second-price auction. The intuition is that bids are lower and the constraint is less likely to bind in a first-price auction than in a second-price auction. Their work is related to our paper because illiquidity also imposes a constraint on feasible bids. The difference is that, in the case of illiquidity, such a constraint depends on the assignment of trade intents to solvers and hence is relevant only in a combinatorial problem. [^4] Dobzinski et al. (2012) study multiple simultaneous auctions with financially constrained bidders. In their paper, however, there is a single budget constraint over the sum of the possible bids, which is quite different from the feasibility constraint generated by illiquidity. We are also related to the literature studying multiple simultaneous auctions, such as, for example, Milgrom (2000), Krishna and Rosenthal (1996), Goeree and Lien (2014), and Gentry et al. (2019). In particular, Gentry et al. (2019) studies simultaneous sealed bit auctions with complementarities, corresponding to one of the benchmarks we consider (modulo the feasibility constraint we introduce).
To the best of our knowledge, the only other papers studying trade intent auctions are Chitra et al. (2024) and Bachu et al. (2024). Chitra et al. (2024) consider a single Dutch auction and study how the equilibrium of the game changes with the number of bidders. Here, instead, we fix the number of bidders and study different mechanisms when complementarities between trades are present, and the mechanism designer has fairness considerations. Bachu et al. (2024) study empirically two implementations of Dutch auction for trade intents (that of 1inch fusion and UniswapX) and find that this mechanism generates between 4 and 5 basis points improvement over Uniswap routing algorithm.
Finally, our paper relates to the literature studying how frictions in the ability to reallocate utility between players affect the equilibrium of cooperative games. In this literature, games in which it is possible to frictionlessly reallocate utility between players are called TU (Transferable Utility) games, and those in which this frictionless reallocation is not possible are called NTU (Non-TU) games. For example, it is well known that several cooperative game theory concepts are equivalent under TU but different under NTU. Also, the equilibrium in matching games with TU may change under NTU (Becker, 1973, Legros and Newman, 2007). The main difference is that, in cooperative games, the reference for fairness (i.e., the outside options) is common knowledge. In contrast, here, this reference needs to be elicited as part of the mechanism.
inlineinlinetodo: inline “While interest in the issues represented by the non-transferable case is both longstanding and lively (see for instance Farrell-Scotchmer, 1988 on production in partnerships; Rosenzweig-Stark, 1989 on risk sharing in households; and more recently, Lazear, 2000 on incentive schemes for workers; Ackerberg-Botticini, 2002 on sharecropping; and Chiappori-Salanié, 2003 on the empirics of contracts)”,
2 The model
There are two traders and two solvers:
Traders.
The first trader wants to sell one unit of asset $A$ for asset $B$. The second trader wants to sell one unit of asset $C$ for asset $D$. The traders’ payoffs are simply the assets they receive: if trader 1 receives $x$ units of asset $B$ and trader 2 receives $y$ units of asset $D$, their payoffs are $u_{1}=x$ and $u_{2}=y$, respectively.
Solvers.
Each solver can produce a given amount of assets as a function of the orders it is matched with. To parsimoniously capture both the benefit of specialization and the additional efficiencies from batching, we assume that:
  • if solver 1 is matched exclusively with order one, it produces $\beta>0$ units of asset B, where $\beta$ is drawn at the beginning of the game from a continuous and atomless distribution over the interval $[\underline{\beta},\overline{\beta}]$ for $0<\underline{\beta}<\overline{\beta}$. If, instead, solver 2 is matched exclusively with order 1, it produces $\underline{\beta}$ units of asset $B$.
  • If solver 2 is matched exclusively with order two, it produces $\delta>0$ units of asset D, where $\delta$ is drawn at the beginning of the game from a continuous and atomless distribution over the interval $[\underline{\delta},\overline{\delta}]$ for $0<\underline{\delta}<\overline{\delta}$. If, instead, solver 1 is matched exclusively with order 2, it produces $\underline{\delta}$ units of asset $D$.
  • If solver 1 is matched with both orders, it produces $g\cdot\beta$ units of asset B and $g\cdot\underline{\delta}$ units of asset D for $g\geq 1$. If solver 2 is matched with both orders, it produces $g\cdot\underline{\beta}$ units of asset B and $g\cdot\delta$ units of asset D. We assume $g\geq 1$ to capture the additional efficiencies when a solver executes both orders.
Hence, each solver is specialized in a given order: solver 1 is better at order 1, and solver 2 is better at order 2. Note also that there is a single way to execute both orders: when matched with both orders, a solver cannot choose to generate more of a given asset at the expense of the other asset (of course, as we’ll see later, when bidding a solver can return more of a given asset at the expense of the other asset). In Section 6.2, we relax this assumption by introducing the choice of how to execute both orders. Finally, the parameters, $\underline{\beta}$, $\overline{\beta}$, $\underline{\delta}$, $\overline{\delta}$, and $g$ are common knowledge among the players but are unknown to the designer, who therefore cannot condition the mechanism on their values (see below).
Denote a matching of trades and solvers by a tuple
$\mu\in\underbrace{\{\overbrace{1,2}^{\mbox{solver}}\}}_{\mbox{Order 1}}\times\underbrace{\{\overbrace{1,2}^{\mbox{solver}}\}}_{\mbox{Order 2}}$
where the first entry corresponds to the solver matched with the first order and similarly for the second entry. Call $x_{1}(B)$ and $x_{1}(D)$ the amount of token $B$ and $D$ returned by solver 1, and similarly define $x_{2}(B)$ and $x_{2}(D)$. Our main assumption is that solvers cannot return to the traders more than the sum of the assets than what they produce and their inventory: [^5] Note the similarity between the feasibility constraint and a budget constraint (already mentioned in the literature review when discussing Che and Gale, 1998): they both impose a constraint on the bids that can be paid. However, in Assumption 1, such constraint is a function of $\mu$; that is, it is contingent on the trades the solver wins.
Assumption 1 (Feasibility constraint).
Call $\omega(B)\geq 0$ solver 1’s inventory of asset $B$, and $\omega(D)\geq 0$ solver 2’s inventory of asset $D$, assumed private information of each solver. It must be that:
  • if $\mu=\{1,1\}$ then $x_{1}(B)\leq g\cdot\beta+\omega(B)$, $x_{1}(D)\leq g\cdot\underline{\delta}$, $x_{2}(B)=0$, $x_{2}(D)\leq\omega(D)$.
  • if $\mu=\{2,2\}$ then $x_{2}(B)\leq g\cdot\underline{\beta}$, $x_{2}(D)\leq g\cdot\delta+\omega(D)$, $x_{1}(B)\leq\omega(B)$, $x_{1}(D)=0$.
  • if $\mu=\{1,2\}$ then $x_{1}(B)\leq\beta+\omega(B)$, $x_{2}(D)\leq\delta+\omega(D)$, $x_{2}(B)=x_{1}(D)=0$,
  • if $\mu=\{2,1\}$ then $x_{2}(B)\leq\underline{\beta}$, $x_{1}(D)\leq\underline{\delta}$, $x_{1}(B)\leq\omega(B)$ $x_{2}(D)\leq\omega(D)$,
Note that, from a solver’s perspective, the mechanism’s outcome is stochastic as it depends on the actions of the other solver. The feasibility constraint should hold in each possible realization of the mechanism. Also, the feasibility constraint restricts the solvers’ ability to be market makers: each solver may use its inventory to fill the order in which it has a competitive advantage. However, no solver has an inventory of both B and D assets and hence, no solver can be a market maker on the B/D pair: if this was possible, solvers could frictionlessly reallocate utility between the traders when trades are small but not when they are large (where “small” and “large” depends on the size of the inventories). Assuming no market making on the B/D pair allows us to focus on the most interesting case.
We can now write the solvers’ payoff as a function of $\mu$, $x_{1}(B)$, $x_{1}(D)$, $x_{2}(B)$, and $x_{2}(D)$ such that the feasibility constraint hold.
Assumption 2.
Call $p_{DB}$ the notional market price of asset $D$ in terms of asset $B$, assumed common knowledge. For $x_{1}(B)$, $x_{1}(D)$ such that Assumption 1 holds, solver 1’s payoff is
$\begin{cases}g\cdot\beta-x_{1}(B)+p_{DB}(g\cdot\underline{\delta}-x_{1}(D))&\mbox{if }\mu=\{1,1\}\\ \beta-x_{1}(B)&\mbox{if }\mu=\{1,2\}\\ -x_{1}(B)+p_{DB}(\underline{\delta}-x_{1}(D))&\mbox{if }\mu=\{2,1\}\\ -x_{1}(B)&\mbox{if }\mu=\{2,2\}.\\ \end{cases}$
Similarly, for $x_{2}(B)$, $x_{2}(D)$ such that Assumption 1 holds, solver 2’s payoff is
$\begin{cases}-p_{DB}\cdot x_{2}(D)&\mbox{if }\mu=\{1,1\}\\ p_{DB}(\delta-x_{2}(D))&\mbox{if }\mu=\{1,2\}\\ \underline{\beta}-x_{2}(B)-p_{DB}\cdot x_{2}(D)&\mbox{if }\mu=\{2,1\}\\ g\cdot\underline{\beta}-x_{2}(B)-p_{DB}(g\cdot\delta-x_{2}(D))&\mbox{if }\mu=\{2,2\}\end{cases}$
Hence, each solver values asset $D$ using the notional market price, even if exchanging asset $D$ for asset $B$ within the auction is impossible. This assumption is justified by timing: if a solver needs to deliver assets to the traders, it will need to do so immediately. However, solvers can exchange the assets in their inventory over a longer period, reducing (here, for simplicity, eliminating) all frictions and fees.
2.1 Timing and mechanism
The timing of the game is the following:
  1. 1.
    First, $\beta$ and $\delta$ are drawn.
  2. 2.
    Then, each solver $i\in\{1,2\}$ sends three bids. Two bids are bids on individual orders, denoted by $q_{i}(B)$ and $q_{i}(D)$, in units of asset $B$ and $D$, respectively. The third bid is a tuple $Q_{i}=\{Q_{i}(B),Q_{i}(D)\}$, interpreted as a bid for the entire batch, that is, a bid for both orders, where $Q_{i}(B)$ is in units of asset $B$ and $Q_{i}(D)$ is in units of asset $D$.
  3. 3.
    Finally, the mechanism determines the matching and the assets each trader should receive.
Formally, define a message profile by
$m\equiv\{q_{1}(B),q_{1}(D),Q_{1}(B),Q_{1}(D),q_{2}(B),q_{2}(D),Q_{2}(B),Q_{2}(D)\}\in\mathbb{R}^{8}_{+}.$
A mechanism is $\Gamma=\left\{f,b_{1},b_{2},d_{1},d_{2}\right\}$ where
  • $f:\mathbb{R}^{8}_{+}\rightarrow\{1,2\}\times\{1,2\}$ determines the matching of trades to solvers,
  • $b_{i}:\mathbb{R}^{8}_{+}\rightarrow\mathbb{R}_{+}$ determines the amount of asset $B$ that solver $i\in\{1,2\}$ should deliver to trader 1,
  • $d_{i}:\mathbb{R}^{8}_{+}\rightarrow\mathbb{R}_{+}$ determines the amount of asset $D$ that solver $i\in\{1,2\}$ should deliver to trader 2,
The mechanism is common knowledge at the beginning of the game.
An important observation is that the feasibility constraint also has implications for the set of feasible mechanisms. Intuitively, a mechanism is not feasible if, for all non-trivial bids by a solver $i\in\{1,2\}$, there are inventories for solver $i$ and bids by the other solvers such that solver $i$ violates the feasibility constraint. Conversely, a mechanism is feasible if, for any possible inventory, solvers can make non-trivial bids and be guaranteed to respect the feasibility constraint.
Definition 1 (Feasible mechanism).
A mechanism $\Gamma$ is feasible if $\forall i\{1,2\}$,
$\exists\{q_{i}(B),q_{i}(D),Q_{i}(B),Q_{i}(D)\}\geq\{\underline{\beta},\underline{\delta},\underline{\beta},\underline{\delta}\}$
with at least one strict inequality, such that
$\forall\{q_{-i}(B),q_{-i}(D),Q_{-i}(B),Q_{-i}(D)\}$
by solver $-i\in\{1,2\}/i$, $b_{i}(m)$ and $d_{i}(m)$ satisfy Assummption 1 for $\omega(B)=\omega(D)=0$ and $\mu=f(m)$.
Hence, the feasibility constraint restricts the mechanisms that can be implemented. For example, for a mechanism to be feasible, it must be the case that if the outcome is specialization, then each winning solver only delivers assets to one trader: for a message profile $m\in\mathbb{R}^{8}_{+}$ such that $f(m)=\{i,-i\}$ we have $b_{-i}(m)=d_{i}(m)=0$. Similarly, if a solver wins no trades, it does not deliver any asset: for $m\in\mathbb{R}^{8}_{+}$ such that $i\notin f(m)$, then $b_{i}(m)=d_{i}(m)=0$. Feasibility, therefore, rules out several well-known mechanisms. Any “all pay” auction mechanism is not feasible. Interestingly, also Vickrey-Clark-Groves (VCG) mechanisms are not feasible. In a VCG mechanism, each winner pays the highest-valued combination of bids when his or her bids are removed. If the equilibrium is specialization, then if one of the winner’s bids is removed, the outcome will be batching. However, this implies that a solver who won only one order and produced only one asset should also deliver the other asset, which is not feasible.
Finally, in the remainder of the paper, we will often say that “the outcome of the auction is determined by the batched bid” or “is determined by the individual bids.” Intuitively, the outcome is determined by the batched bid if the matching of trades to solvers does not change when the individual bids are made worse. Similarly, the outcome is determined by the individual bids if the matching does not change when the batched bid is made worse. The next definition formalizes this intuition.
Definition 2.
For a given mechanism $\Gamma$ and a given message profile
$m=\{q_{1}(B),q_{1}(D),Q_{1}(B),Q_{1}(D),q_{2}(B),q_{2}(D),Q_{2}(B),Q_{2}(D)\}\}$
we say that the outcome is determined by the individual bids if $f(m)=f(m^{\prime})$ for all
$m^{\prime}=\{q_{1}(B),q_{1}(D),Q^{\prime}_{1}(B),Q^{\prime}_{1}(D),q_{2}(B),q_{2}(D),Q^{\prime}_{2}(B),Q^{\prime}_{2}(D)\}$
with $\{Q^{\prime}_{1}(B),Q^{\prime}_{1}(D),Q^{\prime}_{2}(B),Q^{\prime}_{2}(D)\}\leq\{Q_{1}(B),Q_{1}(D),Q_{2}(B),Q_{2}(D)\}$. Similarly, we say that the outcome is determined by the batched bid if $f(m)=f(m^{\prime\prime})$ for all
$m^{\prime\prime}=\{q^{\prime\prime}_{1}(B),q^{\prime\prime}_{1}(D),Q_{1}(B),Q_{1}(D),q^{\prime\prime}_{2}(B),q^{\prime\prime}_{2}(D),Q_{2}(B),Q_{2}(D)\}$
with $\{q^{\prime\prime}_{1}(B),q^{\prime\prime}_{1}(D),q^{\prime\prime}_{2}(B),q^{\prime\prime}_{2}(D)\}\leq\{q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D)\}$.
Note that whether the auction outcome is determined by the batched bid or individual bids differs from whether the outcome is batching or specialization: for a given mechanism, the auction’s outcome may be determined by the individual bids; nonetheless, a single solver may win both orders.
3 Batch auction and simultaneous standard auctions
We can use the above framework to study the two dominant mechanisms for trade intent auctions: simultaneous standard auctions (as implemented by 1inch Fusion and UniswapX in the form of Dutch auctions) and batch auctions (as implemented by CoW Swap). These mechanisms are quantitatively important: in May 2024, 1-inch fusion and UniswapX intermediated USD 650 million and USD 1.6 billion, respectively, while CoW Swap intermediated USD 2.4 billion. These mechanisms are also relevant theoretical benchmarks because they are sub-cases of the above model: simultaneous standard auctions are equivalent to the above mechanism in which the batch bids are ignored, while batched auctions are equivalent to the above mechanism in which individual bids are ignored.
3.1 Simultaneous standard auctions
When both trades are auctioned off simultaneously, solvers anticipate that they may win both trades and generate additional efficiencies. This is related to what the literature calls “exposure problem” arising in multiple simultaneous auctions with complementarities: a bidder who expects to win an auction with strictly positive probability values winning other auctions more and hence may increase its bid accordingly. By doing so, however, the bidder incurs a loss when he or she wins only one auction (see Meng and Gunay, 2017).
In our case, the feasibility constraint limits the amount of “exposure” (i.e., negative profits in the realization): if there is a strictly positive probability that the outcome is specialization, then solver $1$ cannot bid higher than $\underline{\delta}$ for the second order, and solver $2$ cannot bid higher than $\underline{\beta}$ for the first order. Solver 1 will then bid for order 1 knowing that the opponent will bid at most $\underline{\beta}$, and solver 2 will bid on order 2 knowing that the opponent will bid at most $\underline{\delta}$, leading to the equilibrium described in the following lemma (its proof is omitted).
Lemma 1.
If there are two simultaneous standard auctions, then in equilibrium, solver 1 wins the first order and returns $\underline{\beta}$ units of asset $B$, while solver 2 wins the second order and returns $\underline{\delta}$ units of asset $D$ [^6] Note that, technically, there could be other equilibria because the least productive solver on each trade never wins and hence could bid any amount that satisfies its feasibility constraint. If the auction is, for example, in second price, then some equilibria are even worse for the traders than that described in the Lemma. These equilibria are, however, not robust to perturbations of the type space. For example, if solver 1 can produce $\underline{\delta}+\epsilon$ for $\epsilon\rightarrow 0$ when it wins only trade 2, and solver 2 can produce $\underline{\beta}+\epsilon$ when it wins only trade 1, then there an infinitesimally small probability that solver 1 wins trade 2 and solver 2 wins trade 1. Then, the equilibrium described in the Lemma becomes unique.
The above result is quite transparent when the two auctions are in second price because, in each auction, the lowest bid determines the assets returned to the trader by the winning solver. But it extends to any standard auction format. In particular, it holds when the two auctions are in first price, which is strategically equivalent to the Dutch auctions implemented by 1inch Fusion and UniswapX. [^7] In this case, however, there are some technicalities since the bid space is continuous. If the space of bids is discrete and the two auctions are in first price, then in equilibrium the “strong” solver in each auction will outbid the “weak” solver by the minimal amount. If the bid space is continuous, instead, in the unique equilibrium, the two solvers bid the same amount, and then an endogenous tie-breaking rule as in Jackson et al. (2002) is used to choose the winner (which is always the strongest solver). See also Gentry et al. (2019), who use the endogenous tie-breaking rule to show the existence of the equilibrium in simultaneous standard auctions with complementarities.
Traders, therefore, receive the minimum amount of assets. Intuitively, specialization means that the auction is not competitive: there is a “strong,” and a “weak” solver in each auction, and the strong solver can win by matching the offer of the “weak” solvers. The productivity of the weak solver in each auction then determines the assets that traders receive.
3.2 Batch auction
To start, remember that feasibility prevents “second price” mechanisms that use the losing bid to determine the assets returned by the winning solver: even if the winning bid has a higher value, it may nonetheless deliver less of one asset than the second-highest bid. We, therefore, restrict our attention to a “first price” (or “pay your bid”) mechanism in which the winning solver returns the assets specified in its batched bid, which is the approach followed by CoW Swap. [^8] In practice, the batch auction model implemented by CoW Swap also has a subsidy to solvers, which we ignore for simplicity.
In a batch auction, each bid is evaluated based on the market value of the assets returned to the traders. There is, therefore, a degree of indeterminacy, as the same market value can be generated via different bids by varying the amounts of asset $B$ and asset $D$ returned to the traders. For the moment, we consider a specific equilibrium: one in which each solver bids the maximum on the order in which it has a competitive disadvantage. In such an equilibrium, $Q^{*}_{2}(B)=g\underline{\beta}$ and $Q^{*}_{1}(D)=g\underline{\delta}$. Define $V\equiv g\underline{\beta}-p_{DB}\cdot g\underline{\delta}$ and write solver 1 best response as
$\mbox{argmax}_{Q_{1}(B)\leq g\beta}\left\{\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V\right\}\right\}$
and solver 2’s best response as
$\mbox{argmax}_{Q_{2}(D)\leq g\delta}\left\{p_{DB}\left(\delta\cdot g-Q_{2}(D)\right)\mbox{pr}\left\{p_{DB}Q_{2}(D)+V>Q_{1}(B)\right\}\right\}.$
We now introduce the following change of variables: $x=Q_{1}(B)+V$, and $y=p_{DB}Q_{2}(D)$. By doing so, we can rewrite the solvers’ best responses as
$\mbox{argmax}_{x}\left\{(g\beta-V-x)\mbox{pr}\{x\geq y\}\right\}$
$\mbox{argmax}_{y}\left\{(p_{DB}\cdot g\delta-y)\mbox{pr}\{y\geq x\}\right\},$
which are the two best responses of a first-price auction with types’ distributed over $[g\underline{\beta}-V,g\overline{\beta}-V]$ and $[p_{DB}\cdot g\underline{\delta},p_{DB}g\overline{\delta}]$, leading to the following lemma (its proof is omitted).
Lemma 2.
There is an equilibrium of the batch auction with bids $Q^{*}_{2}(B)=g\underline{\beta}$, $Q^{*}_{1}(D)=g\underline{\delta}$ and $Q^{*}_{1}(B),Q^{*}_{2}(D)$, where $\{Q^{*}_{1}(B)+V,\leavevmode\nobreak\ p_{DB}Q^{*}_{2}(D)\}$ are the equilibrium bids of a first price auction with types distributed over $[g\underline{\beta}-V,g\overline{\beta}-V]$ and $[p_{DB}\cdot g\underline{\delta},p_{DB}g\overline{\delta}]$ (appropriately derived from the underlying distribution of the parameters $\beta$ and $\delta$).
However, as already mentioned, solvers care only about the total value of their bid, which implies that the equilibrium is not unique. That is, given the equilibrium bids derived in the above lemma, any $\tilde{Q}_{1}(B)\in[0,g\cdot\beta]$ and $\tilde{Q}_{1}(D)\in[0,g\cdot\underline{\delta}]$ such that
$\tilde{Q}_{1}(B)+p_{DB}\tilde{Q}_{1}(D)=Q^{*}_{1}(B)+p_{DB}Q^{*}_{1}(D)>g\underline{\beta}+p_{DB}g\underline{\delta}$
are also equilibrium bids by solver 1. Similarly, any $\tilde{Q}_{2}(B)\in[0,g\cdot\underline{\beta}]$ and $\tilde{Q}_{2}(D)\in[0,g\cdot\delta]$ such that
$\tilde{Q}_{2}(B)+p_{DB}\tilde{Q}_{2}(D)=Q^{*}_{2}(B)+p_{DB}Q^{*}_{2}(D)>g\underline{\beta}+p_{DB}g\underline{\delta}$
are also equilibrium bids by solver 2.
The important observation is that, in the batch auction, the total value of the assets returned to the traders always exceeds that of the simultaneous individual-trade auctions. The reason is that batching forces the two solvers to compete: even if there is a single “strong” solver per trade (and little competition if trades are auctioned off independently), there are two “strong” solvers in the batch auctions, and hence higher competition between solvers. At the same time, there is an indeterminacy relative to what each trader receives. There are equilibria in which a trader receives less in a batch auction than what this trader would have received in the simultaneous individual-trade auctions. We summarize these observations in the following remark.
Remark 1.
There are equilibria of the batch auction in which traders receive more than $\underline{\beta}$ and $\underline{\delta}$, and both traders are better off relative to the simultaneous standard auctions (c.f, Lemma 1). There are also equilibria in which either trader 1 receives less than $\underline{\beta}$ or trader 2 receives less than $\underline{\delta}$, that is, there are equilibria in which one of the two traders receives less in a batch auction than in the simultaneous standard auctions.
Hence, the batch auction has equilibria that, intuitively, seem unfair. To formalize this intuition, the next section introduces a notion of fairness applicable to combinatorial auctions.
4 Fair combinatorial auctions.
The goal of this section is to formalize a notion of fairness applicable to a generic combinatorial auction $\Gamma$. Doing so requires establishing a counterfactual mechanism against which to compare the outcome of $\Gamma$. The key observation is that such counterfactual is part of the mechanism $\Gamma$, because it depends on the bids and determines the equilibrium allocation and payments. To say it differently, the reference for fairness must be generated endogenously as part of the mechanism.
We establish this counterfactual in two steps. First, we restrict the space of mechanisms $\Gamma$ to those in which a form of independence holds: the auction outcome is independent of the batched bid whenever the individual bids determine the outcome. We then note that, under this restriction, if the batched bid determines the auction’s outcome, there is a unique natural counterfactual: a mechanism that is identical to $\Gamma$ but for the fact that the batched bids are ignored.
Assumption 3.
The auction’s outcome is determined either by the individual bids or the batched bids. Furthermore, for all message profiles $m$ such that the individual bids determine the outcome, we have:
$b_{i}(m)=\tilde{b}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))$
$d_{i}(m)=\tilde{d}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))$
for some functions $\tilde{b}_{i}(),\tilde{d}_{i}():\mathbb{R}^{4}\rightarrow\mathbb{R}$.
The first part of the above assumption imposes a mild form of monotonicity in the auction’s allocation. The second part is a form of independence of the auction’s payments whenever the individual bids determine the outcome. It is slightly more restrictive than the first part but still quite mild, especially in conjunction with the requirement that the mechanism is feasible.
Remember that feasibility requires that a solver who does not win trade 1 does not need to return any asset B and, similarly, a solver who does not win trade 2 does not need to return any asset D. It also implies that if a solver wins a trade, there is an upper bound to the assets that such a solver needs to return. This upper bound is solely a function of the winner’s bids so that, when bidding, a solver can be sure not to violate the feasibility constraint for any bids by the opponent. If, for simplicity, we interpret this upper bound as the bid on the individual trade itself, then feasibility implies that the winner of trade 1 (or 2) returns an amount of assets B (or D) that is between zero and its bid on trade 1 (or 2). The above assumption then further implies that, in case the outcome is determined by the individual bids, where exactly the winner’s payment lies in the feasible interval is independent of the batched bids.
We are now ready to introduce our notion of fairness.
Definition 3 (Fairness).
A mechanism is fair relative to a given $\{\tilde{b}(.,.,.,),\tilde{d}(.,.,.,.)\}$ if and only if
$\sum_{i=1}^{2}b_{i}(m)\geq\sum_{i=1}^{2}\tilde{b}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))$
and
$\sum_{i=1}^{2}d_{i}(m)\geq\sum_{i=1}^{2}\tilde{d}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))$
Note that $\{\tilde{b}(.,.,.,),\tilde{d}(.,.,.,.)\}$ are themselves mechanisms. For example, they could be two simultaneous first-price auctions:
$\begin{split}\tilde{b}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))=\begin{cases}q_{i}(B)&\mbox{ if }q_{i}(B)>q_{-i}(B)\\ 0&\mbox{ otherwise }\end{cases}\\ \tilde{d}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))=\begin{cases}q_{i}(D)&\mbox{ if }q_{i}(D)>q_{-i}(D)\\ 0&\mbox{ otherwise }\end{cases}\end{split}$ (1)
or two simultaneous second-price auctions:
$\begin{split}\tilde{b}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))=\begin{cases}q_{-i}(B)&\mbox{ if }q_{i}(B)>q_{-i}(B)\\ 0&\mbox{ otherwise }\end{cases}\\ \tilde{d}_{i}(q_{1}(B),q_{1}(D),q_{2}(B),q_{2}(D))=\begin{cases}q_{-i}(D)&\mbox{ if }q_{i}(D)>q_{-i}(D)\\ 0&\mbox{ otherwise }\end{cases}\end{split}$ (2)
Importantly, even if (1) or (2) hold, the reference for fairness is not the equilibrium of those auctions (as derived in Lemma 1). The reason is that, as we will see later in more detail, solvers may bid differently on individual trades if they know that their bids will be used to determine the reference for fairness.
Also important to note is that every mechanism is trivially fair for some $\{\tilde{b}(.,.,.,),\tilde{d}(.,.,.,.)\}$. For example, any auction in which the batched bids never determine the outcome (such as the simultaneous standard auction) is trivially fair. Also, the batch auction is fair whenever $\tilde{b}(.,.,.,)=\tilde{d}(.,.,.,.)\equiv 0$. In a sense, every mechanism is fair to itself. The above definition bites whenever we compare two different mechanisms. For example, in the next section, we will consider a mechanism that allows batching only when it is fair relative to (1) or (2).
5 Equilibrium of the fair combinatorial auction
The fair combinatorial auction introduces several novel strategic considerations. First, each solver may bid higher in individual auctions to disqualify the opponent’s batched bid as unfair. Similarly, a solver may submit a higher batched bid to avoid being disqualified as unfair. At the same time, if a solver expects the opponent’s batched bid to be disregarded as unfair, then this solver should submit a lower batched bid. Finally, also in this case, the feasibility constraint limits the amount that each solver can bid. As we will see, the way these different strategic considerations shape the equilibrium is mediated by the auction format.
The remainder of the section, therefore, derives the equilibrium of different types of fair combinatorial auctions. For ease of comparison with the above benchmarks, we restrict our attention to mechanisms such that:
  • if the individual bids determine the outcome of the combinatorial auction, it is equivalent to two simultaneous first-price auctions as in (1), or two simultaneous second-price auctions as in (2).
  • for all $m$ such that the outcome of the combinatorial auction is determined by the batched bids, then the winning solver pays its bids: whenever $f(m)=\{i,i\}$, then
    $b_{i}(m)=Q_{i}(B)\leavevmode\nobreak\ \leavevmode\nobreak\ d_{i}(m)=Q_{i}(D)\leavevmode\nobreak\ \leavevmode\nobreak\ b_{-i}(B)=b_{-i}(D)=0.$ (3)
    If multiple batched bids are fair, then the winner is the one with higher market value. [^9] This assumption is both for simplicity and for consistency with the standard batch auction. Also, the intuition behind our results extends to other monotonically increasing functions of the assets returned to the users.
5.1 Second-price individual auctions.
Suppose that the individual, item-by-item auctions are second-price auctions. Formally, define the solver with the best batched-bid as $W\equiv\mbox{argmax}_{i\in\{1,2\}}\left\{Q_{i}(B)+p_{DB}Q_{i}(D)\right\}$. Then, the mechanism is:
  • if $Q_{W}(B)>\min\{q_{1}(B),q_{2}(B)\}$ and $Q_{W}(D)>\min\{q_{1}(D),q_{2}(D)\}$, then $f(m)=\{W,W\}$, $b_{W}(m)=Q_{W}(B)$, $d_{W}(m)=Q_{W}(D)$, and all other transfers are zero. [^10] Also here, there are some technicalities related to how ties are broken in case a solver matches the opponent’s batched bid with its individual-trade bid. The formulation in the text assumes that, in case of ties, the batched bid is disqualified as unfair. This assumption guarantees the existence of the equilibrium; otherwise, a solver may want to outbid the opponent’s batched bid by the smallest possible amount, but because the action space is continuous, this “smallest possible amount” is not defined. We also note that an alternative to imposing such a tie-breaking rule is to introduce an endogenous tie-breaking rule as in Jackson et al. (2002), which also guarantees that the equilibrium when the action space is discrete converges to the equilibrium with continuous action space as the action space becomes finer (see also the discussion in footnote7).
  • Otherwise, the best bid on a given order wins it and pays the second-highest bid. Formally, denote the winner of order 1 by $w_{1}\equiv\mbox{argmax}_{i\in\{1,2\}}\{q_{i}(B)\}$, and the winner of order 2 by $w_{2}\equiv\mbox{argmax}_{i\in\{1,2\}}\{q_{i}(D)\}$. The mechanism in this case is: $f(m)=\{w_{1},w_{2}\}$, $b_{w_{1}}(m)=\min\{q_{1}(B),q_{2}(B)\}$, $d_{w_{2}}(m)=\min\{q_{1}(D),q_{2}(D)\}$ and all other transfers are zero.
Consider any strictly positive $Q_{1}(B),Q_{2}(B),Q_{1}(D),Q_{2}(D)$ that are an equilibrium of the batch auction discussed earlier. The key observation is that these bidding strategies are also an equilibrium of the fair combinatorial auction with individual bids in second price whenever $q_{1}(B)=q_{2}(B)<\min\{Q_{1}(B),Q_{2}(B)\}$ and $q_{1}(D)=q_{2}(D)<\min\{Q_{1}(D),Q_{2}(D)\}$. The reason is that no individual solver can manipulate the reference for fairness. That is, it is possible that, in equilibrium, $q_{1}(B)=q_{2}(B)<Q_{i}(B)<\underline{\beta}$ $\forall i\{1,2\}$; nonetheless, no solver can change its individual-trade bid and disqualify the opponent’s bid as unfair. Finally, the above argument holds for any $Q_{1}(B),Q_{2}(B),Q_{1}(D),Q_{2}(D)$ strictly positive, while it fails (due to our tie-breaking rule) whenever the winning bid of the batch auction delivers zero assets to a trader, leading to the following lemma.
Lemma 3.
The set of equilibria of the fair combinatorial auction with second-price individual trade auctions (as in 2) is the set of equilibria of the batch auction in which each trader receives strictly positive amounts.
Intuitively, when the individual-trade auctions are in second price, each individual solver has limited ability to manipulate the reference for fairness. Depending on the players’ bids, it is even possible that the outcome of the auction gives almost no asset to one player, and nonetheless, no solver can disqualify such an outcome as unfair. As a consequence, this auction reduces to, again, the batch auction (c.f., Lemma 2), with the only difference that delivering exactly zero assets to a trader is not an equilibrium here (but it is an equilibrium of the batch auction).
Finally, note how, in this case, the mechanism used to define fairness is the second-price auction, but a trader may receive less than what they would receive in two simultaneous second-price auctions. That is, in equilibrium, the fair combinatorial auction does not guarantee that each trader receives at least as much as they would have received in the equilibrium of the mechanism used as a reference for fairness. This is not a general property: in the next mechanism we study, the fair combinatorial auction guarantees that each trader receives strictly more than what they would have received in the equilibrium of the mechanism used as a reference for fairness.
5.2 First-price individual auctions.
The mechanism, in this case, is:
  • if $Q_{W}(B)>\max\{q_{1}(B),q_{2}(B)\}$ and $Q_{W}(D)>\max\{q_{1}(D),q_{2}(D)\}$, then $f(m)=\{W,W\}$, $b_{W}(m)=Q_{W}(B)$, $d_{W}(m)=Q_{W}(D)$, and all other transfers are zero.
  • Otherwise, $f(m)=\{w_{1},w_{2}\}$, $b_{w_{1}}(m)=q_{w_{1}}(B)$, $d_{w_{2}}(m)=q_{w_{2}}(D)$ and all other transfers are zero.
We start with a useful preliminary result.
Lemma 4.
Suppose that, in equilibrium:
$\begin{split}Q^{*}_{1}(B)+p_{DB}Q^{*}_{2}(D)\geq g\underline{\beta}+p_{DB}g\underline{\delta}Q^{*}_{2}(B)+p_{DB}Q^{*}_{2}(D)\geq g\underline{\beta}+p_{DB}g\underline{\delta},\end{split}$ (4)
that is, each solver’s batched bid returns at least the minimum value produced. Then, it must be that $Q^{*}_{1}(D)=g\cdot\underline{\delta},\leavevmode\nobreak\ Q^{*}_{1}(B)\in[g\underline{\beta},g\beta)$, and similarly $Q^{*}_{2}(B)=g\underline{\beta},\leavevmode\nobreak\ Q^{*}_{2}(D)\in[g\underline{\delta},g\cdot\delta)$.
The lemma shows that, in its batched bid, each solver bids the maximum on the trade in which it has a disadvantage. The reason is that a solver’s batched bid is more likely to be considered unfair due to the amount of assets returned on the “weak” order (which is the opponent’s “strong” order). On the other component of the batched bid, each solver delivers more than the minimum amount but less than the maximum amount.
{NiceTabular}
c||c|c|[hvlines] & $q_{1}(B)<g\underline{\beta}$ $q_{1}(B)\geq g\underline{\beta}$
[…]
$q_{2}(D)\geq g\underline{\delta}$ (2,2) $(1,2)$
Table 1: Matching of trades to solvers as a function of the bids not pinned down by Lemma 4.
An immediate corollary of the above lemma is that, in equilibrium, a solver can disqualify the opponent’s batched bid only with the individual bid on the trade in which it is stronger. Using this result, Table 5.2 plots the auction outcome as a function of the solvers’ bids. We can, therefore, write solver $1$’s payoff as:
$\begin{cases}(\beta-q_{1}(B))\mbox{pr}\left\{q_{2}(D)\geq g\underline{\delta}\right\}+\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}&\mbox{ if }q_{1}(B)\geq g\underline{\beta}\\ \left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}&\mbox{ otherwise }\end{cases}$
where, again, $V\equiv g\underline{\beta}-p_{DB}\cdot g\underline{\delta}$. Note that, as long as $q_{1}(B)\geq g\underline{\beta}$, whether solver 1 wins one or both trades is independent of its bids, and solver 1 should set $q_{1}(B)=Q_{1}(B)=g\underline{\beta}$. Using this, we can rewrite solver 1’s payoff as
$\begin{cases}(\beta-g\underline{\beta})\mbox{pr}\left\{q_{2}(D)\geq g\underline{\delta}\right\}+\left(\beta\cdot g-g\underline{\beta}\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}&\mbox{ if }q_{1}(B)=g\underline{\beta}\\ \left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}&\mbox{ otherwise }\end{cases}$
A second observation is that, as long as (4) holds, then $Q_{1}(B)\geq g\underline{\beta}$ and
$\left(\beta\cdot g-g\underline{\beta}\right)$ $\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}>$
$\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{q_{2}(D)<g\underline{\delta}\right\}\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}$
and solver 1 always prefers $q_{1}(B)=g\underline{\beta}$ (i.e., kicking out the opponent’s batched bid) to $q_{1}(B)<g\underline{\beta}$ (i.e., not kicking out the opponent’s batched bid). However, $q_{1}(B)=g\underline{\beta}$ may not be feasible. Finally, note that when $q_{1}(B)<g\underline{\beta}$, then solver 1 payoff does not depend on $q_{1}(B)$ and, without loss of generality, we can assume $q_{1}(B)=\underline{\beta}$.
It follows that solver 1 best response is:
$\begin{cases}q_{1}(B)=Q_{1}(B)=g\underline{\beta}&\mbox{ if }\beta\geq g\underline{\beta}\\ q_{1}(B)=\underline{\beta},\leavevmode\nobreak\ Q_{1}(B)=\tilde{Q}_{1}(B)&\mbox{ otherwise }\end{cases}$
where
$\tilde{Q}_{1}(B)\equiv\mbox{argmax}_{Q_{1}(B)}\left\{\left(\beta\cdot g-Q_{1}(B)\right)\mbox{pr}\left\{Q_{1}(B)>p_{DB}Q_{2}(D)+V|q_{2}(D)<g\underline{\delta}\right\}\right\}$
Repeating the same reasoning for solver 2 yields the best response:
$\begin{cases}q_{2}(D)=Q_{2}(D)=g\underline{\delta}&\mbox{ if }\delta\geq g\underline{\delta}\\ q_{2}(D)=\underline{\delta},\leavevmode\nobreak\ Q_{2}(D)=\tilde{Q}_{2}(D)&\mbox{ otherwise }\end{cases}$
where
$\tilde{Q}_{2}(D)\equiv\mbox{argmax}_{Q_{2}(D)}\left\{p_{DB}\left(\delta\cdot g-Q_{2}(D)\right)\mbox{pr}\left\{p_{DB}Q_{2}(D)+V>Q_{1}(B)|q_{1}(B)<g\underline{\beta}\right\}\right\}$
$\beta$$\delta$$\underline{\beta}$$g\cdot\underline{\beta}$$\overline{\beta}$$\underline{\delta}$$g\cdot\underline{\delta}$$\overline{\delta}$$q_{1}(B)=Q_{1}(B)=g\underline{\beta}$$q_{2}(D)=Q_{2}(D)=g\underline{\delta}$(specialization)$q_{1}(B)=Q_{1}(B)=g\underline{\beta}$$q_{2}(D)=\underline{\delta};\leavevmode\nobreak\ Q_{2}(D)=\tilde{Q}_{2}(D)$(batching solver 1)$q_{1}(B)=\underline{\beta};\leavevmode\nobreak\ Q_{1}(B)=\tilde{Q}_{1}(B)$$q_{2}(D)=Q_{2}(D)=g\underline{\delta}$(batching solver 2)$q_{1}(B)=\underline{\beta};\leavevmode\nobreak\ Q_{1}(B)=\tilde{Q}_{1}(B)$$q_{2}(D)=\underline{\delta};\leavevmode\nobreak\ Q_{2}(D)=\tilde{Q}_{2}(D)$(competitive batching)
Figure 1: The two best responses
Figure 1 illustrates the two best responses in the $(\beta,\delta)$ space. It highlights the four possible regimes that may emerge in equilibrium, which we discuss in the next proposition.
Proposition 1.
Suppose the individual trade auctions are in first price as in (1). Then, in equilibrium, (4) holds and:
  • If $\beta<g\underline{\beta}$ and $\delta<g\underline{\delta}$, then $q^{*}_{1}(B)=\underline{\beta}$, $q^{*}_{2}(D)=\underline{\delta}$, and $\{Q^{*}_{1}(B)-V,p_{DB}Q^{*}_{2}(D)\}$ are the equilibrium bids in a first price auction with types distributed over $[g\underline{\beta}-V,\min\{g^{2}\underline{\beta}-V,g\overline{\beta}-V\}]$ and $[p_{DB}\cdot g\underline{\delta},p_{DB}\cdot\min\{g^{2}\underline{\delta},g\overline{\delta}\}]$ (with distributions appropriately derived from the underlying distribution of the parameters $\beta$ and $\delta$). The winning solver wins both trades and the equilibrium is batching,
  • If $\beta\geq g\underline{\beta}$ and $\delta\geq g\underline{\delta}$, then solver 1 bids $q^{*}_{1}(B)=Q^{*}_{1}(B)=g\underline{\beta}$, solver 2 bids $q^{*}_{2}(D)=Q^{*}_{2}(D)=g\underline{\delta}$, solver 1 wins trade 1 and solver 2 wins trade 2,
  • If $\beta\geq g\underline{\beta}$ and $\delta<g\underline{\delta}$, then solver 1 bids $q^{*}_{1}(B)=Q^{*}_{1}(B)=g\underline{\beta}$ and solver 2 bids $q^{*}_{2}(D)=\underline{\delta},\leavevmode\nobreak\ Q^{*}_{2}(D)=\tilde{Q}_{2}(D)$. Solver 1 wins both trades.
  • If $\beta<g\underline{\beta}$ and $\delta\geq g\underline{\delta}$, then solver 1 bids $q^{*}_{1}(B)=\underline{\beta},\leavevmode\nobreak\ Q^{*}_{1}(B)=\tilde{Q}_{1}(B)$, and solver 2 bids $q^{*}_{2}(D)=Q^{*}_{2}(D)=g\underline{\delta}$. Solver 2 wins both trades.
The remaining equilibrium bids are derived in Lemma 4.
Proof.
The only missing part of the proof is to show that, in the first case, $\tilde{Q}_{1}(B)$ and $\tilde{Q}_{2}(D)$ are best responses of a first price auction. To do so, it is enough to follow the same steps discussed in Section 3.2: define $x=Q_{1}(B)+V$ and $y=p_{DB}Q_{2}(D)$, and rewrite the best responses. ∎
The first case corresponds to “competitive batching”. The main difference with the batch auction (see Lemma 2) is that here the equilibrium is unique: the traders receive $g\underline{\beta}$ units of asset B and $g\underline{\delta}$ units of asset D. Concerning the value of the assets returned to the users, if $g\underline{\delta}>\overline{\delta}$ and $g\underline{\beta}>\overline{\beta}$ the equilibrium is identical to that of the batch auction. If, instead, $\beta<g\underline{\beta}$ and $\delta<g\underline{\delta}$ but either $g\underline{\delta}<\overline{\delta}$ or $g\underline{\beta}<\overline{\beta}$, then the equilibrium may be different from that of the batched auction because, here, the type distribution is truncated.
The second case corresponds to specialization. The equilibrium allocation is identical to the simultaneous standard auctions, but traders receive more assets (c.f. Lemma 1). The reason is that each solver uses its individual bid to disqualify the opponent’s batched bid as unfair: if either solver were to drop its bid below the equilibrium level, the opponent’s batched bid would win.
Finally, the remaining two cases correspond to “uncompetitive batching”: the benefit of specialization is in some intermediate range, and one of the solvers is very productive (relative to the other). This solver can easily disqualify the opponent’s batched bid as unfair. Anticipating no competition from the opponent’s batched bid, this solver wins by bidding the minimal amount to win both orders.
To summarize, when the individual auctions are in first price, each solver can more directly affect the reference for fairness and, as a consequence, will bid higher in the individual auctions. The auction, therefore, provides stronger fairness guarantees because the two traders earn more than $g\underline{\beta}$ and $g\underline{\delta}$, against $\underline{\beta}$ and $\underline{\delta}$ in the simultaneous first-price auctions (c.f., Lemma 1) and in the fair combinatorial auction with individual auctions in second price (c.f., Lemma 3). At the same time, the value of the assets returned to the traders will be lower, especially if the outcome is specialization or “uncompetitive batching”. The comparison with the previous mechanism highlights a trade off in the design of fair combinatorial auctions: providing stronger fairness guarantees comes at the expense of lowering the value of the assets returned to the traders.
6 Discussion.
We now discuss a few ways in which the basic model can be extended.
6.1 Sequential fair combinatorial auction
A drawback of the above mechanism is that solvers submit batched bids without knowing the reference for fairness. It is, therefore, possible that a solver’s batched bid is disqualified as unfair while this solver could have produced a fair batched bid (for example, if the individual trade auctions are in first price and the outcome is specialization).
An intuitive solution is to have a sequential combinatorial auction: first, solvers bid on the individual trades; then these bids (and the reference outcome for fairness) are revealed; then solvers submit batched bids. The sequential combinatorial auction, however, generates information leakage: how a solver bids in the first phase reveals some information to the opponent, who may use it in the subsequent bidding stage. Because of this, in equilibrium, solvers submit only “uninformative” bids, which is equivalent to skipping the first bidding stage and only bidding on batches, as the next Proposition shows.
Proposition 2.
In every pure strategy equilibrium of the sequential fair combinatorial auction, in the first stage, solver 1 bids fixed amounts $\hat{q}_{1}(B)<g\underline{\beta}$ and $\hat{q}_{1}(D)<g\underline{\delta}$ independent of the realization of $\beta$. Similarly, in the first stage, solver 2 bids fixed amounts $\hat{q}_{2}(B)<g\underline{\beta}$ and $\hat{q}_{2}(D)<g\underline{\delta}$ independent of the realization of $\delta$. The second stage of the auction is identical to a batch auction where each solver delivers strictly more than $\hat{q}_{1}(B)$ and $\hat{q}_{2}(D)$.
Note that because $\hat{q}_{1}(B)$ and $\hat{q}_{2}(D)$ are undetermined (i.e., there are multiple equilibria), the set of equilibria is identical to that of the simple batch auction.
6.2 “Unfair” trades executions
One limitation of the above model is that unfair outcomes occur exclusively because there is an indeterminacy regarding how solvers bid in the batch auction. Hence, in a batch auction, next to “unfair” equilibria, there are always also “fair” equilibria. This is an artifact of the model’s simplicity: in this section, we show that if solvers have multiple ways to execute the two orders, then all equilibria of the batch auction may be “unfair”.
To see this, assume that a solver who wins both orders can execute them in two ways. If solver 1 wins both orders, it can choose to generate either $\{g\beta,g\underline{\delta}\}$ (like before) or $\{k\underline{\beta},\tau\underline{\delta}\}$ with $\tau<1$ and $1<g<k$. If, instead, solver 2 wins both orders, it can choose to generate either $\{g\underline{\beta},g\delta\}$ (like before) or $\{k\underline{\beta},\tau\underline{\delta}\}$. We further assume that $k\underline{\beta}>g\overline{\beta}$, so the second option (i.e., $\{k\underline{\beta},\tau\underline{\delta}\}$) always generates more of asset $B$ but less of asset $D$ relative to the other option.
Figure 2 provides a graphical representation of the technology available to Solver 1 (on the right panel) and Solver 2 (on the left panel), together with the total value of each possible execution of both trades. The important observation is that, as $p_{BD}$ changes, the execution that generates the highest value in case a solver wins both orders also changes. In particular, if $p_{BD}$ is sufficiently low, then for both solvers, $\{k\underline{\beta},\tau\underline{\delta}\}$ generates a higher value than the alternatives. Finally, the figure also illustrates the space of feasible bids in case a solver wins both orders, as a function of the choice of production (where we assume zero inventory for simplicity). Note that when the choice of production is $\{k\underline{\beta},\tau\underline{\delta}\}$, then feasibility implies that trader 2 must receive less than $\underline{\delta}$.
Asset $B$Asset $D$$\tau\cdot\underline{\delta}$$\underline{\delta}$$g\cdot\underline{\delta}$$\underline{\beta}$$\beta$$g\cdot\beta$$k\cdot\underline{\beta}$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\beta+p_{DB}\cdot\tau\cdot\underline{\delta}$
Asset $B$Asset $D$$\tau\cdot\underline{\delta}$$\underline{\delta}$$\delta$$g\cdot\delta$$\underline{\beta}$$g\cdot\underline{\beta}$$k\cdot\underline{\beta}$$g\cdot\beta+p_{DB}\cdot g\cdot\underline{\delta}$$k\cdot\underline{\beta}+p_{DB}\cdot\tau\cdot\underline{\delta}$
Figure 2: Solvers’ production possibilities when matched with both orders (red and green dots) and when matched with a single order (black dot). The shaded areas are the sets of feasible transfers in case a solver wins both orders, as a function of the choice of production.
It turns out that this model variation delivers results identical to those presented earlier. The only difference is that depending on the parameters, in all equilibria of the batch auction, one trader receives less than in the simultaneous standard auctions, even though it is possible to provide both traders with strictly more assets. The equilibrium of the simultaneous standard auctions is unchanged: trader 1 receives $\underline{\beta}$ and trader 2 receives $\underline{\delta}$ (see Lemma 1). The equilibrium of the batch auction is similar to that derived in the main text, but with one difference: if $p_{DB}$ is sufficiently low, then both solvers will produce $\{k\underline{\beta},\tau\underline{\delta}\}$ when matched with both orders. In this case, competition implies that trader 1 receives $k\beta$ and trader 2 receives $\tau\underline{\delta}$. As a consequence, trader 2 receives less than $\underline{\delta}$ in all equilibria of the game, even though it is possible to produce and return tokens in a way that makes both traders better off relative to the two simultaneous standard auctions. If $p_{DB}$ is sufficiently low, the unique equilibrium of the batch auction is, intuitively, unfair.
Concerning the fair combinatorial auction, when the individual-trade auction is in second price, again, the logic discussed earlier continues to hold: there is an equilibrium in which both traders set the individual-trade bids to zero, and the auction does not provide any additional fairness guarantee relative to the batch auction. Finally, when the individual-trade auction is in first price, then the equilibrium is again identical to the one derived earlier: solvers never produce $\{k\underline{\beta},\tau\underline{\delta}\}$ because, if they were to do so, their batched bit would always be disqualified as unfair. Again, both traders earn at least $g\underline{\beta}$ and $g\underline{\delta}$, and the auction is effective at providing some minimal fairness guarantees.
inlineinlinetodo: inlineThe fact that batch auctions are always better than two simultaneous auctions is a strong result. As am extension, we may want to check whether it holds also in a more general version of the model (in the sense of having two arbitrarily type distributions).
6.3 Generalization to multiple traders: fairness as a filter
A natural question is how to extend the above framework to more than two traders and more than two solvers. One way to do so is to have a standard combinatorial auction in which solvers submit multiple batched bids (in addition to individual-trade bids), and then some of these batched bids are filtered out because unfair.
Such an extension is relevant because fairness has implications for the computational complexity of the combinatorial auction. It is well known that, as the number of trades increases, determining the combination of bids that maximizes the value of the tokens returned can be computationally intractable (sometimes called the winner determination problem). At the same time, the fairness filter also changes. If, as the number of trades increases, the fairness filter becomes more stringent, then the computational complexity of the problem may remain bounded. Formally exploring the computational complexity of fair combinatorial auctions and how it changes with the number of trades is left for future work.
7 Conclusion
We study trade-intent auctions. These auctions are quantitatively relevant, as they intermediate approximately USD 5 billion monthly. They are also theoretically interesting. There are production complementarities between trades, and the design is, therefore, combinatorial. However, unlike other combinatorial auctions studied in the literature, trade-intent auctions lack a common numeraire that can be used to reallocate the efficiencies of batching among the different traders.
We started by building a theoretical model of trade intent auctions with possible complementarities between trade intents. We then used this model to study the two most common auction formats: (i) batch auctions, in which different trade intents are auctioned off jointly as a batch; and (ii) simultaneous standard auctions, in which there is a separate auction for each trade intent (of which Dutch auctions are a subcase). The main result of this comparison is that batch auctions generate more competition between solvers than simultaneous standard auctions. The reason is that batching forces all solvers to compete: even in cases where there is a “best” solver at each trade, the batch auction forces each “best” solver to outbid the other to win any trade. At least under the assumptions of our model, we find that batch auctions always deliver higher overall value to the traders than individual trade-intent auctions (such as Dutch auctions). At the same time, in a batch auction, the solvers’ payoff only depends on the value of the assets returned to the traders, not how this value is shared. Therefore, the benefit of the batch auction may accrue disproportionally to one trader, leaving the other worse off relative to the simultaneous standard auctions.
We then use the model to study a new type of auction: the fair combinatorial auction. This type of auction can be reduced to a combination of batch auctions and simultaneous standard auctions, because solvers can submit bids on individual trades and also on batches of trades. The key assumption is that the outcome of the individual trade-intent auctions (as constructed using the individual-trade bids) is used to filter out batched bids: batched bids are considered by the auctioneer only when they are better than the outcome of the individual trade auctions for all traders in the batch. The outcome of the individual trade auctions can, therefore, be seen as a benchmark for fairness: all batched solutions that improve upon this outcome for all traders are “fair” and considered in the auction, and all other batched solutions are discarded as “unfair”.
The key observation is that, even if the mechanism is a combination of batch auctions and simultaneous standard auctions, there is no guarantee that its equilibrium is a combination of the equilibrium of the batch auctions and that of the individual trade-intent auctions. The reason is that solvers anticipate that their bid on the individual trades will be used to construct the reference for fairness. Hence, they may bid differently on individual trades in the fair combinatorial auction than in the simultaneous standard auctions (with no batching possible). In designing such an auction, it is crucial to properly account for these novel strategic incentives, as they determine the fairness guarantees that the auction provides in equilibrium.
We illustrates this point by showing that when the individual trade-intent auctions are second-price auctions, the fair combinatorial auction does not provide any additional fairness guarantees relative to the simple batch auction. The reason is that there are equilibria in which all solvers bid the minimum amount on individual trades, which is equivalent to skipping the individual auctions and only having a batch auction. If the individual auctions are in second price, then no solver can change this outcome by changing its bid on the individual auctions. Instead, when the individual trade-intent auctions are first-price auctions, each solver can manipulate the reference for fairness with its bids on individual trades, independently of the bids by other solvers. In this case, we show that the fair combinatorial auction guarantees that all traders receive more than what they would earn in the simultaneous standard auctions. Hence, even though not all fair combinatorial auctions provide strong fairness guarantees in equilibrium, some do. However, these higher fairness guarantees come at the expense of lower value returned to the traders.
Mathematical Appendix
Proof of Lemma 4.
To show that $Q^{*}_{2}(B)=g\underline{\beta}$ and $Q^{*}_{1}(D)=g\underline{\delta}$, we proceed by contradiction: suppose that $Q^{*}_{2}(B)<g\underline{\beta}$. The fact that the total value of the batched bid must be greater or equal to $g\underline{\beta}+p_{BD}g\underline{\delta}$ implies that $Q^{*}_{2}(D)>g\delta$. Hence, the probability that solver 2’s batched bid is considered unfair because of order 2 is zero, while this probability is strictly positive for order 1. Solver 2 is better off by decreasing $Q^{*}_{2}(D)$ and increasing $Q^{*}_{2}(B)$ to keep the total value of the batched bid constant because doing so reduces the probability that the batched bid is considered as unfair. Hence, it must be that $Q^{*}_{2}(B)=g\cdot\underline{\beta},\leavevmode\nobreak\ Q^{*}_{2}(D)<g\delta$. The same reasoning can be applied to show that $Q^{*}_{1}(B)<g\beta,\leavevmode\nobreak\ Q^{*}_{1}(D)=g\cdot\underline{\delta}$.
Proof of Proposition 2.
First, we prove that, in any equilibrium, solvers bid fixed amounts in the first bidding stage. We do so by contradiction. Suppose that, in equilibrium, solver 1’s first stage bid is an increasing function of $\beta$ (note: it could be globally increasing or locally increasing). If this is the case, upon observing a first-period bid from solver 1, solver 2 can infer the maximum batched bid that solver 1 can submit and try to outbid solver 1 by a minimal amount. Anticipating this, in the first period of the game, solver 1 should always bid the minimal amount to deceive solver 2 into submitting a low batched bid. Hence, solver 1’s submitting a first-stage bid is an increasing function of $\beta$ is not an equilibrium. An identical argument rules out equilibria in which solver 2’s submits a first-stage bid that is an increasing function of $\delta$. inlineinlinetodo: inlinethe proof shows that there cannot be a pure strategy equilibrium, but I believe the argument extends to mixed strategy equilibria: in these equilibria, the realized bid does not exactly reveal the type, but does allow the opponent to make some inference and hence adjust the batched bid
To close the proposition, note that if all first-stage bids are below $g\underline{\beta}$ and $g\underline{\delta}$, then with probability 1 there is batching in equilibrium. The first-stage bids do not matter, and feasibility is never violated. Conversely, if a first-stage bid is greater than $g\underline{\beta}$ or $g\underline{\delta}$, there are values of $\beta$ and $\delta$ for which the feasibility constraint is violated. Such bids would, therefore, be informative of the realization of $\beta$ and $\delta$, which cannot be an equilibrium. ∎